The complex number z satisfies the equation 2z-iz* = 3(3-5i) where z* denotes the complex conjugate of z. Determine the value of z giving your answer in the form x+yi where x and y are real numbers.
1
Expert's answer
2020-11-05T17:03:24-0500

Let

"z = x + iy \Rightarrow z* = x - iy"

Then

"\begin{array}{l}
2(x + iy) - i(x - iy) = 3(3 - 5i)\\
2x + 2iy - ix + {i^2}y = 9 - 15i\\
2x - y + i(2y - x) = 9 - 15i
\end{array}"

Then

"\left\{ \begin{array}{l}
2x - y = 9\\
2y - x = - 15
\end{array} \right."

"\left\{ \begin{array}{l}
2x - y = 9\\
3x = 3
\end{array} \right."

"\left\{ \begin{array}{l}
x = 1\\
y = -7
\end{array} \right."

Answer: "z = 1 - 7i"